### v1

# Curvatures and Non-metricities in the Non-Relativistic Limit of Bosonic Supergravity

Abstract
We construct a diffeomorphism-covariant formulation of the non-relativistic (NR) limit of bosonic supergravity. This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to the NR degrees of freedom. The construction is purely geometrical and is based on a torsionless connection. The non-metricities are associated with the gravitational fields of the theory, $\tau_{\mu\nu},h_{\mu\nu},\tau^{\mu\nu}$ and $h^{\mu\nu}$, and are fixed by requiring compatibility with the relativistic metric. We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature. Our results establish an equivalence between the proposed construction and the intrinsic torsion framework of string Newton–Cartan geometry. We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
I Introduction
In recent years, non-relativistic (NR) limits of string theory NR1 -NR3 and supergravity NRST1 -NRST13 have attracted considerable attention, motivated by both conceptual and more phenomenological developments. In particular, the systematic understanding of the supergravity limit of NR string theory has led to the emergence of Newton–Cartan and string Newton–Cartan geometries as the natural geometric frameworks underlying these limits (for reviews and complementary introductions, see Review -Review3 and references therein). Let us start by considering the universal bosonic NS–NS supergravity, whose relativistic field content consists of a spacetime metric $\hat{g}_{\mu\nu}$, a Kalb–Ramond two-form $\hat{B}_{\mu\nu}$ and a dilaton $\hat{\phi}$. A controlled NR limit is obtained NSNS by introducing an explicit expansion in powers of the speed of light $c$,
$\hat{g}_{\mu\nu}$ $=$ $c^{2}\tau_{\mu\nu}+h_{\mu\nu}\,$, (1)
$\hat{g}^{\mu\nu}$ $=$ $\frac{1}{c^{2}}\tau^{\mu\nu}+h^{\mu\nu}\,$, (2)
$\hat{B}_{\mu\nu}$ $=$ $-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,$, (3)
$\hat{\phi}$ $=$ $\ln(c)+\varphi\,$, (4)
where $\mu,\nu=0,\dots,25$ are space-time indices and the longitudinal directions are labeled by $a,b=0,1$. We also have
$\tau_{\mu\nu}$ $=$ $\tau_{\mu}{}^{a}\tau_{\nu}{}^{b}\eta_{ab}\,$, (5)
$\tau^{\mu\nu}$ $=$ $\tau^{\mu}{}_{a}\tau^{\nu}{}_{b}\eta_{ab}\,$, (6)
$h_{\mu\nu}$ $=$ $e_{\mu}{}^{a^{\prime}}e_{\nu}{}^{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (7)
$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (8)
according to the vielbein decomposition of the fields ($a^{\prime}=2,...,25$). Moreover, they satisfy the defining relations,
$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,$,
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
Although individual geometric quantities constructed from the relativistic Levi–Civita connection exhibit divergences in the limit $c\to\infty$, the bosonic NS–NS action,
$S=\int d^{26}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big)\,$, (10)
with $\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}$, remains finite due to a non-trivial cancellation between divergent contributions arising from the Ricci scalar and the Kalb–Ramond sector. This observation has been central to the development of other NR string supergravity theories such as heterotic formulations BandR - Sigmaheterotic , supersymmetric extensions NRST11 -NRST12 and Double Field Theory DFT1 -DFT4 , in the context of non-Riemannian double geometries NRDFT1 -NRDFT9 and its NR limit EandD .
Motivated by the progress on the NR limit of string supergravities, different approaches to study higher-derivative corrections under the NR limit have recently emerged Higher-Derivative1 -Higher-Derivative4 . Some of the theories already explored can be written in terms of tensors with respect to infinitesimal diffeomorphisms in their relativistic level. In order to be able to decompose these tensors preserving the covariance, one needs to construct a suitable connection with respect to the longitudinal and transverse parts of the metric degrees of freedom (and its inverse), and rewrite these contributions in terms of curvatures and covariant derivatives with respect to this symmetry. The same can be done for the Lorentz inherited symmetries, like boosts, $SO(1,1)$ and $SO(8)$ transformations, which was done for the NR limit of bosonic supergravity in NSNS at the two-derivative level. In particular, a fully diffeomorphism-covariant formulation of this theory, formulated directly in terms of intrinsic NR geometric data and their associated curvatures is still lacking. The main obstruction lies in the fact that the relativistic Levi–Civita connection does not admit a smooth NR limit when expressed solely in terms of Newton–Cartan variables.
In this work, we solve this issue by constructing a torsionless affine connection that is explicitly covariant under spacetime diffeomorphisms. Crucially, this connection is not metric compatible with respect to the fundamental NR fields. Instead, the covariant derivatives of $\tau_{\mu}{}^{a}$ and $h^{\mu\nu}$ are controlled by specific non-metricity tensors, which are uniquely fixed by requiring consistency with the relativistic metric condition. As a result, the full bosonic supergravity action can be rewritten in a manifestly covariant form with respect to diffeomorphisms, where all terms remain finite and well defined as $c\rightarrow\infty$.
The purpose of this paper is threefold. First, we present an explicit construction of a torsionless connection and the associated non-metricities within the framework of Newton–Cartan gravity. Second, we compute the corresponding curvatures and show explicitly how all relativistic curvature tensors reorganize into NR geometric invariants. Third, we discuss several applications of the formalism. These include the derivation of the complete non-relativistic bosonic supergravity Lagrangian in a fully covariant form, the systematic analysis of higher-derivative contributions such as Riemann-squared and other four-derivative bosonic corrections, and the construction of more general Newton–Cartan–inspired theories with alternative choices of non-metricity that relax boost invariance beyond the bosonic supergravity setting.
Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity. We expect this framework to serve as a useful starting point for higher-derivative extensions and for exploring duality-covariant formulations of string theory in the NR limit.
II Infinitesimal diffeomorphism transformations and covariant derivatives
We start by recalling the transformation rule of the fundamental fields with respect to infinitesimal diffeomorphisms,
$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau_{\mu}{}^{a}}+\partial_{\mu}{\xi^{\rho}}\tau_{\rho}{}^{a}\,$,
$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau^{\mu}{}_{a}}-\partial_{\rho}{\xi^{\mu}}\tau^{\rho}{}_{a}\,$,
$\delta_{\xi}h_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h_{\mu\nu}}+2\partial_{(\mu}{\xi^{\rho}}h_{\nu)\rho}\,$,
$\delta_{\xi}h^{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h^{\mu\nu}}–2\partial_{\rho}{\xi^{(\mu}}h^{\nu)\rho}\,.$ (11)
The torsionless connection is given by
$\Gamma_{\mu\nu}^{\rho}(\tau,h)=\frac{1}{2}h^{\rho\sigma}(2\partial_{(\mu}h_{\nu)\sigma}-\partial_{\sigma}h_{\mu\nu})$
[…]
which is the $c^{0}$-contribution coming from the relativistic connection, preserving the non-tensorial transformation rule, in the sense that it is the contribution which transforms as a connection.
The covariant derivatives are therefore defined as
$\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}\,$, (13)
$\nabla_{\mu}\tau^{\nu}{}_{a}$ $=$ $\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\sigma}^{\nu}\tau^{\sigma}{}_{a}\,$, (14)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $\partial_{\mu}h_{\nu\rho}-2\Gamma_{\mu(\nu}^{\sigma}h_{\rho)\sigma}\,$, (15)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $\partial_{\mu}h^{\nu\rho}+2\Gamma_{\mu\sigma}^{(\nu}h^{\rho)\sigma}\,.$ (16)
The Riemann tensor can be easily computed from the commutator of the covariant derivatives,
$[\nabla_{\mu},\nabla_{\nu}]\tau^{\rho}{}_{a}=R^{\rho}{}_{\epsilon\mu\nu}\tau^{\epsilon}{}_{a}\,$, (17)
giving the usual expression
$R^{\rho}{}_{\epsilon\mu\nu}(\tau,h)=2\partial_{[\mu}\Gamma_{\nu]\epsilon}^{\rho}+2\Gamma_{[\mu|\alpha}^{\rho}\Gamma_{|\nu]\epsilon}^{\alpha}\,.$ (18)
III Non-metricities
The non-metricities play a fundamental role in the theory, since they guarantee that the NR limit can be taken from the relativistic theory,
$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $Q^{(\tau)}_{\mu\nu\rho}\,$, (19)
$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $Q^{(\tau^{-1})}_{\mu}{}^{\nu\rho}\,$, (20)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $Q^{(h)}_{\mu\nu\rho}\,$, (21)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $Q^{h^{-1}}_{\mu}{}^{\nu\rho}\,$. (22)
The fixed values of the non-metricities come from the metric conditions,
$\hat{\nabla}_{\mu}g_{\nu\rho}$ $=$ $0\,$,
$\hat{\nabla}_{\mu}g^{\nu\rho}$ $=$ $0\,$,
where $\hat{\nabla}$ is the usual covariant derivative with respect the relativistic Levi–Civita connection. Therefore, after doing the NR expansion, one needs the following non-metricity combinations,
$Q^{(\tau)}_{\mu\nu\rho}$ $=$ $e^{\alpha a^{\prime}}(\partial_{\mu}\tau_{(\nu\alpha}+\partial_{(\nu}\tau_{\mu\alpha}-\partial_{\alpha}\tau_{\mu(\nu})e_{\rho)a^{\prime}}$ (23)
$Q^{(\tau^{-1})}_{\rho}{}^{\mu\nu}$ $=$ $-\tau^{(\mu\alpha}(2\partial_{(\rho}h_{\sigma)\alpha}-\partial_{\alpha}h_{\rho\sigma})h^{\sigma\nu)}$ (24)
$Q^{(h)}_{\mu\nu\rho}$ $=$ $\tau^{\alpha}{}_{b}(\partial_{\mu}h_{(\nu\alpha}+\partial_{(\nu}h_{\mu\alpha}-\partial_{\alpha}h_{\mu(\nu})\tau_{\rho)}{}^{b}$ (25)
$Q^{h^{-1}}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
The previous quantities are not boost invariants  [^1] Here $\lambda_{aa^{\prime}}$ is a generic boost parameter NSNS .,
$\delta_{\lambda}h_{\mu\nu}$ $=$ $-2\lambda_{aa^{\prime}}e_{(\mu}{}^{a^{\prime}}\tau_{\nu)}{}^{a}\,$, (27)
$\delta_{\lambda}\tau^{\mu\nu}$ $=$ $2\lambda_{a}{}^{a^{\prime}}e^{(\mu}{}_{a^{\prime}}\tau^{\nu)a}\,$, (28)
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-2\epsilon_{ab}\lambda^{a}{}_{a^{\prime}}\tau_{[\mu}{}^{b}e_{\nu]}{}^{a^{\prime}}\,$. (29)
and consequently the non-metricities cannot be set to zero without compromising the boost symmetry in bosonic supergravity. We will return to this point in the discussion section, since one might be interested in exploring Newton-Cartan geometries with more arbitrary non-metricities, beyond bosonic supergravity.
IV Decomposition of the relativistic curvatures
IV.1 The Riemann tensor
The relativistic Riemann tensor is given by
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{(4)}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{(2)}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
$+c^{(-2)}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{(-4)}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
So far we have just identified part of the covariant contribution inside $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$, given by $R^{\rho}{}_{\epsilon\mu\nu}$ but other covariant contributions are encoded in $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ due to $c^{2}$- and $c^{-2}$-contributions in the relativistic Levi-Civita connection. All the contributions in the relativistic Riemann tensor can be written in covariant form, and this is what we are going to show now. Let us start with the higher-order contribution,
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\nu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
$(\partial_{\nu]}\tau_{\epsilon\beta}+\partial_{\epsilon}\tau_{\nu]\beta}-\partial_{\beta}\tau_{\nu]\epsilon})\,.$
The previous quantity transforms as a tensor, and it can be written in a manifest covariant form using the covariant derivatives,
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
$(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
The same technique can be implemented for the lowest contribution of the Riemann tensor,
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\nu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
$(\partial_{\nu]}h_{\epsilon\beta}+\partial_{\epsilon}h_{\nu]\beta}-\partial_{\beta}h_{\nu]\epsilon})$
which therefore can be written as
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
$(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$
In these cases we have identified that the structure of $R^{(-4)\rho}{}_{\epsilon\mu\nu}$ is just an interchange of $\tau_{\mu\nu}\leftrightarrow h_{\mu\nu}$ and $h^{\mu\nu}\leftrightarrow\tau^{\mu\nu}$ with respect to the highest order contribution, $R^{(4)\rho}{}_{\epsilon\mu\nu}$.
The $c^{2}$-contributions of the relativistic Riemann tensor takes an important role, since it is part of bosonic supergravity action once one takes the NR limit. These contributions can be written in terms of the covariant derivatives
$\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}=2\nabla_{[\mu}\Big[h^{\rho\alpha}(\nabla_{\nu]}\tau_{\epsilon\alpha}+\nabla_{\epsilon}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\epsilon})\Big]\,.$ (35)
Again, we can interchange $\tau_{\mu\nu}\leftrightarrow h_{\mu\nu}$ and $h^{\mu\nu}\leftrightarrow\tau^{\mu\nu}$ to construct the $c^{(-2)}$-contributions,
$\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}=2\nabla_{[\mu}\Big[\tau^{\rho\alpha}(\nabla_{\nu]}h_{\epsilon\alpha}+\nabla_{\epsilon}h_{\nu]\alpha}-\nabla_{\alpha}h_{\nu]\epsilon})\Big]\,.$ (36)
Finally, we proceed to construct the $c^{0}$-contributions to the relativistic Riemann tensor. This quantity is given by,
$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
In the next section we will study the decomposition of the (relativistic) Ricci tensor in terms of the Newton-Cartan fields, and in covariant form with respect to infinitesimal diffeomorphisms.
IV.2 The Ricci tensor
The relativistic Ricci tensor is given by
$\hat{R}_{\epsilon\nu}$ $=$ $c^{4}\hat{R}^{(4)}_{\epsilon\nu}+c^{2}\hat{R}^{(2)}_{\epsilon\nu}+\hat{R}^{(0)}_{\epsilon\nu}+c^{-2}\hat{R}^{(-2)}_{\epsilon\nu}+c^{-4}\hat{R}^{(-4)}_{\epsilon\nu}\,$. (38)
Using the results of the previous subsection, we can directly write all the contributions in covariant form,
$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
$\hat{R}^{(2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[h^{\mu\alpha}(\nabla_{\nu]}\tau_{\epsilon\alpha}+\nabla_{\epsilon}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\epsilon})\Big]\,$, (40)
$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
$\hat{R}^{(-2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[\tau^{\mu\alpha}(\nabla_{\nu]}h_{\epsilon\alpha}+\nabla_{\epsilon}h_{\nu]\alpha}-\nabla_{\alpha}h_{\nu]\epsilon})\Big]\,.$ (42)
$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
IV.3 The Ricci scalar
The relativistic Ricci scalar is decomposed as
$\hat{R}$ $=$ $c^{2}\hat{R}^{(2)}+\hat{R}^{(0)}+c^{-2}\hat{R}^{(-2)}+c^{-4}\hat{R}^{(-4)}\,$, (44)
where the covariant form of each contribution is given by
$\hat{R}^{(2)}$ $=$ $\hat{R}^{(2)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(4)}_{\mu\nu}\tau^{\mu\nu}\,$, (45)
$\hat{R}^{(0)}$ $=$ $\hat{R}^{(0)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(2)}_{\mu\nu}\tau^{\mu\nu}\,$, (46)
$\hat{R}^{(-2)}$ $=$ $\hat{R}^{(-2)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(0)}_{\mu\nu}\tau^{\mu\nu}\,$, (47)
$\hat{R}^{(-4)}$ $=$ $\hat{R}^{(-4)}_{\mu\nu}h^{\mu\nu}\,.$ (48)
Using the results of the previous subsection we can easily see that the right hand side of the previous contributions can be written in a fully covariant way. We will explore the explicit form of $\hat{R}^{(0)}$ in the next section.
V Applications
In this section, we explore three direct applications of the proposed formalism. First, we derive the finite non-relativistic limit of the two-derivative bosonic supergravity Lagrangian. We then discuss the convenience of this formulation for extracting covariant contributions at any desired order in the expansion in powers of $c$. As an illustrative example, we compute all finite contributions arising from the Metsaev–Tseytlin Lagrangian, which constitutes part of the four-derivative corrections to the bosonic Lagrangian in the NR limit [^2] Additional contributions are expected to arise in order to cancel the higher-derivative divergences of the action.. Finally, we discuss the possibility of employing more general non-metricities, focusing in particular on the case in which all such tensors vanish. This choice explicitly breaks the boost symmetry of the NR limit of bosonic supergravity.
V.1 NR-Bosonic supergravity Lagrangian: covariant (re-)formulation
This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the $\hat{H}-$terms, and using the results of the previous section it can be written in the following way,
$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
$+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}$
$+2\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\nu]}\tau_{\mu\alpha}+\nabla_{\mu}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi$
$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
where
$h_{\mu\nu\rho}$ $=$ $3\partial_{[\mu}b_{\nu\rho]}\,,\quad\hat{H}^{(2)}_{\mu\nu\rho}=-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,$.
Now we can explicitly see that the finite bosonic Lagrangian after taking the NR limit contains a $R_{\mu\nu}h^{\mu\nu}$ contribution, plus other terms related to the non-metricities. The construction is equivalent to the torsionful-formalism of string Newton-Cartan.
V.2 Non-relativistic bosonic $\alpha^{\prime}$-corrections
The four-derivative corrections to the bosonic string supergravity were historically computed considering three- and four-point scattering amplitudes for the massless states GrossSloan -CN (see Electure for a pedagogical review of this topic). The method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
$S_{MT}=\int d^{D}x\sqrt{-g}e^{-2\phi}(L^{(0)}+L^{(1)}_{MT})\,$, (50)
where
$L^{(1)}_{\rm MT}$ $=$ $-\frac{\alpha^{\prime}}{4}\Big[\hat{R}_{\mu\nu\rho\sigma}\hat{R}^{\mu\nu\rho\sigma}-\frac{1}{2}\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}+\frac{1}{24}\hat{H}^{4}-\frac{1}{8}\hat{H}^{2}_{\mu\nu}\hat{H}^{2\mu\nu}\Big]\,$,\, (51)
and
$\hat{H}^{2}_{\mu\nu}$ $=$ $\hat{H}_{\mu}{}^{\rho\sigma}\hat{H}_{\nu\rho\sigma}\,,\quad\hat{H}^{2}=\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\,.$ (52)
The formalism developed in this letter allows to easily inspect the NR contributions of the previous quantities without losing covariance. For example, recalling that $\alpha^{\prime}\rightarrow\frac{\alpha^{\prime}_{NR}}{c^{2}}$ Higher-Derivative3 -Higher-Derivative4 , the finite contributions coming from $\hat{R}_{\mu\nu\rho\sigma}\hat{R}^{\mu\nu\rho\sigma}$ are given by the $c^{2}$-contributions of the previous term, after the NR decomposition,
$2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}h_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+4\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(0)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}h_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}h_{\mu\gamma}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}h_{\mu\gamma}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}\tau^{\sigma\lambda}h^{\nu\epsilon}$
$+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}h_{\mu\gamma}h^{\nu\epsilon}h^{\sigma\lambda}+\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}h_{\mu\gamma}h^{\rho\delta}h^{\sigma\lambda}$
$+\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}\tau^{\sigma\lambda}h^{\nu\epsilon}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}h^{\nu\epsilon}$
$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}$ (53)
Similarly, one can inspect the rest of the terms. For $\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}$ the finite contributions are given by
$\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\gamma\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\sigma\alpha}h_{\mu\gamma\delta}h_{\epsilon\lambda\alpha}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\rho\gamma}h^{\nu\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\epsilon\lambda}h_{\gamma\delta\alpha}$
$+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\sigma}{}_{\gamma\epsilon\delta}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}h^{\nu\epsilon}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\gamma\lambda}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\lambda}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\gamma\lambda}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\mu}{}_{\sigma\gamma\epsilon}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}\tau^{\gamma\lambda}h^{\rho\sigma}h^{\epsilon\alpha}h_{\delta\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}\tau^{\epsilon\lambda}h^{\nu\sigma}h^{\rho\alpha}h_{\delta\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\gamma\delta}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(-2)\mu}{}_{\delta\lambda\alpha}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}h^{\rho\delta}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}h^{\rho\gamma}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}\tau^{\epsilon\alpha}h^{\rho\delta}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}\tau^{\gamma\lambda}h^{\epsilon\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\rho\epsilon}\,.$ (54)
The terms only containing $\hat{H}-$contributions can also be easily written in covariant form. For example, $H^{4}$ contributes with the following finite terms,
$4\hat{H}^{(2)}_{\mu\nu\rho}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\sigma\gamma\epsilon}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+6\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$+6\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+24\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\lambda\alpha}h_{\beta\chi\psi}$
$+12\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\delta\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+36\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\gamma}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$+12\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\nu\chi}h^{\rho\psi}h^{\sigma\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\beta\chi\psi}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\rho\epsilon}h^{\delta\beta}h^{\lambda\chi}h^{\alpha\psi}$
$+18\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\delta\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}\,$, (55)
while $H^{2}_{\mu\nu}H^{2\mu\nu}$ contributes with,
$4\hat{H}^{(2)}_{\mu\nu\rho}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\sigma\gamma\delta}h_{\epsilon\alpha\chi}h_{\lambda\beta\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\sigma}h^{\rho\gamma}h^{\epsilon\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\lambda\chi}h_{\alpha\beta\psi}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\lambda\chi}h_{\alpha\beta\psi}$
$+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\alpha}h^{\rho\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\alpha\beta}h_{\lambda\chi\psi}+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\lambda\beta}h_{\alpha\chi\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\beta\chi}h_{\lambda\alpha\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\delta}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\delta}h^{\rho\gamma}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+16\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\gamma}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\beta}h^{\rho\chi}h^{\gamma\delta}h^{\epsilon\lambda}h^{\alpha\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\nu\chi}h^{\rho\sigma}h^{\gamma\delta}h^{\epsilon\lambda}h^{\alpha\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\beta}h^{\rho\gamma}h^{\epsilon\delta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\lambda}h^{\rho\alpha}h^{\gamma\chi}h^{\epsilon\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\sigma\chi}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\rho\delta}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\delta\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}$ (56)
At the present moment we cannot control all the divergences of the full Metsaev-Tseytlin Lagrangian and for that reason we are not giving an exhaustive analysis of the four-derivative contributions of this Lagrangian in its NR limit. However, the covariant framework developed in this work could be very useful for attacking this open problem.
V.3 General f(R,Q) non-relativistic theories
While in this work we have shown that the non-metricities (23)-(26) need to be fixed in a very particular way to provide the relativistic metric conditions (III)-(III), one can construct more general NR theories with arbitrary non-metricities. One interesting example is to demand the compatibility condition on the Newton-Cartan fields,
$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $0\,$, (57)
$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $0\,$, (58)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $0\,$, (59)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $0\,$, (60)
which automatically implies that both the relativistic Riemann and Ricci tensor are now finite,
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $R^{\rho}{}_{\epsilon\mu\nu}\,$, (61)
$\hat{R}_{\epsilon\nu}$ $=$ $R_{\epsilon\nu}\,$, (62)
and the more general (gravitational) Lagrangian in this setup is given by
$L_{NR}|_{Q=0}=a_{1}R_{\mu\nu}h^{\mu\nu}+a_{2}R_{\mu\nu}\tau^{\mu\nu}\,$, (63)
with $a_{1}$ and $a_{2}$ two arbitrary coefficients. In this case, the $SO(8)$ and $SO(1,1)$ symmetry from the vielbein formalism can be preserved, but the boost symmetry is broken. Further theories with these properties can be explored by changing the non-metricity contributions and including $L(R,Q)$ in the $L_{NR}$ Lagrangian.
VI Discussion
The present construction provides a geometrical formulation of the non-relativistic limit of bosonic supergravity in terms of a Newton–Cartan geometry with non-metricities. Within this framework, an appropriate choice of connection allows spacetime diffeomorphism covariance to remain manifest throughout the construction. As a result, the full two-derivative bosonic supergravity in its NR limit can be written in a covariant form with respect to infinitesimal diffeomorphisms, thereby complementing other approaches based on subgroups of the Lorentz symmetry NSNS .
A key advantage of this formalism is that it allows for a straightforward decomposition of bosonic higher-derivative contributions, such as $\alpha^{\prime}$-corrections. These corrections are naturally formulated in the metric formalism, where the Lorentz symmetry is manifest and no vielbein (or spin connection) is required. From a practical perspective, this provides a systematic method for analyzing higher-derivative terms by expressing relativistic curvature invariants directly in terms of non-relativistic geometric data. In particular, the use of fixed non-metricities enables a transparent identification of divergent and finite contributions, a task that is typically challenging.
Although we do not yet have full control over the divergences that appear at the four-derivative level, the formalism developed in this work already proves useful for isolating and organizing several finite contributions to the four-derivative bosonic Lagrangian in the NR limit. In particular, we have explicitly demonstrated how to compute the contributions arising from the Metsaev–Tseytlin Lagrangian. At present, however, no mechanism is known to cancel the remaining higher-derivative divergences. We expect that the covariant rewriting presented here will provide a valuable starting point for future investigations aimed at resolving these divergences in the bosonic $\alpha^{\prime}$-corrections, as well as for extending the analysis to heterotic supergravity and more general non-relativistic gravity models.
Acknowledgements.
This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.